# Randomization Tests for Conditional Group Symmetry

### v1

@@ D.4 Sampling from approximation versus true conditional distribution
Figure 13: Rejection rates for $\mathrm{SO}(d)$-equivariance tests that reuse a comparison set (top) and that use independent comparison sets (bottom) for varying values of the $Y|X$ covariance matrix off-diagonal values $p$ and sample size $n$. The dotted lines represent tests that sample from the true conditional $P_{\Gamma|[X]}$, and the solid lines represent tests that sample from an approximation.
@@
Figure 14: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests that reuse a comparison set for the approximate versus exact sampling experiment at sample size $n=250$. Parameter $p$ is the covariance between any two dimensions.
Figure 15: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests that use independent comparison sets for the approximate versus exact sampling experiment at sample size $n=250$. Parameter $p$ is the covariance between any two dimensions.
D.5 Sensitivity to non-equivariance in the mean
@@
$Y\mid X\sim\begin{cases}\mathsf{N}(X+s1_{d},\mathbf{I}_{d})&\text{if }\cos^{-1}\left(\frac{X_{1}}{\sqrt{X_{1}^{2}+X_{2}^{2}}}\right)<p\\ \mathsf{N}(X,\mathbf{I}_{d})&\text{otherwise}\end{cases}\;.$
@@
    A proportion $p$ of rotations that leads to a shift in the mean, determined by the 2D angle between the vector spanned by the first two entries of $X$ and the axis of the first dimension. The conditional distribution is equivariant at both $p=0$ and $p=1$, and most non-equivariant at $p=0.5$.
@@
For various combinations of $p$ and $s$ and sample sizes ranging from $25$ to $500$, the test rejection rates are shown in Figure 16. Our tests are sensitive to both parameters, with powers increasing as the conditional distribution moves away from equivariance in either parameter. When a comparison set is reused, the randomization tests outperform the baseline tests in terms of power.
Figure 16: Rejection rates for tests that reuse a comparison set (top) and that use independent comparison sets (bottom) in the non-equivariant mean experiment, controlled by proportion $p$ and shift $s$ parameters, for varying values of sample size $n$. The conditional distribution is not $\mathrm{SO}(3)$-equivariant in all settings shown.
D.6 Permutation equivariance
@@
In this setting, a convenient maximal invariant to condition on is the order statistic $M(X)=(X_{(1)},\dotsc,X_{(d)})$. For $d\in\{3,4,5\}$ and various values of $s$, the test rejection rates are shown in Figure 17. The results are similar to that of the other synthetic experiments, though we note that the powers of the tests rapidly decreases with increasing dimensions in this particular setting.
Figure 17: Rejection rates for tests that reuse a comparison set (top) and that use independent comparison sets (bottom) for the permutation equivariance experiment for varying values of sample size $n$. The conditional distribution is $\mathbb{S}_{d}$-equivariant only at $s=0$.
D.7 Distribution of transverse momenta in LHC dijet data
Figure 18 shows the distribution of transverse momenta for a sample of $500$ first and second leading constituents from the LHC dijet data. Both distributions visually appear to be $\mathrm{SO}(2)$-invariant.
@@ D.7 Distribution of transverse momenta in LHC dijet data
[Acknowledgments] The authors would like to thank Arthur Gretton and Antonin Schrab for a helpful conversation about randomization tests and MMD estimators.

### v2

@@ E.4 Randomization using representative inversions
Figure 11: Rejection rates for $\mathrm{SO}(d)$-equivariance tests for varying values of the $Y|X$ covariance matrix off-diagonal values $c$ and sample size $n$. The dotted lines represent tests that randomize using a deterministic representative inversion, and the solid lines represent tests that randomize using the inversion kernel. (Note the different vertical scale for $c=0$ versus $c>0$.)
@@
Figure 12: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests that use deterministic representative inversions at sample size $n=250$. Parameter $c$ is the covariance between any two dimensions.
E.5 Test runtime
Table 3 shows the mean wall time (in seconds, computed across 1000 simulations) for the FUSE test for $\mathrm{SO}(d)$-equivariance on 100 observations generated the same way as in Section 6.2.1. Observe that the runtime increases linearly with the number of randomizations $B$, and a single run of the test takes no longer than a minute even with 10,000 randomizations. We also see that reusing the comparison set can noticeably decrease computation time, while the benefits of using a deterministic representative inversion are less significant in this setting. [^5] The results suggest that using a deterministic inversion can sometimes slow down the test, but this is likely due to other factors, such as variations across server allocations. It can also be observed that the effect of data dimension on runtime is minimal with our particular implementation for these tests.
Note that the FUSE test generally has a longer runtime compared to that of the baseline and SK tests, which we do not show here.
Table 3: Mean wall time (in seconds) across $1000$ simulations for variations of the FUSE test for $\mathrm{SO}(d)$-equivariance on samples of size $n=100$.
$B$
Reuse comparison set Deterministic inversion $10^{2}$ $10^{3}$ $10^{4}$
$d=3$ $0.32$ $3.29$ $29.66$
$\checkmark$ $0.24$ $2.22$ $22.52$
$\checkmark$ $0.30$ $3.09$ $28.28$
$\checkmark$ $\checkmark$ $0.25$ $2.19$ $24.41$
$d=6$ $0.32$ $3.41$ $34.31$
$\checkmark$ $0.24$ $2.27$ $25.13$
$\checkmark$ $0.32$ $3.04$ $32.91$
$\checkmark$ $\checkmark$ $0.23$ $2.35$ $21.14$
$d=9$ $0.35$ $3.07$ $33.02$
$\checkmark$ $0.25$ $2.34$ $24.97$
$\checkmark$ $0.32$ $3.06$ $31.48$
$\checkmark$ $\checkmark$ $0.24$ $2.27$ $25.10$
E.6 Sensitivity to non-equivariance in the mean
@@
$Y\mid X\sim\begin{cases}\mathsf{N}(X+s1_{d},\mathbf{I}_{d})&\text{if }\cos^{-1}\left(\frac{X_{1}}{\sqrt{X_{1}^{2}+X_{2}^{2}}}\right)<q\\ \mathsf{N}(X,\mathbf{I}_{d})&\text{otherwise}\end{cases}\;.$
@@
    A proportion $q$ of rotations that leads to a shift in the mean, determined by the 2D angle between the vector spanned by the first two entries of $X$ and the axis of the first dimension. The conditional distribution is equivariant at both $q=0$ and $q=1$, and most non-equivariant at $q=0.5$.
@@
For various combinations of $q$ and $s$ and sample sizes ranging from $25$ to $500$, the test rejection rates are shown in Figure 13. Our tests are sensitive to both parameters, with powers increasing as the conditional distribution moves away from equivariance in either parameter.
Figure 13: Rejection rates for tests in the non-equivariant mean experiment, controlled by proportion $q$ and shift $s$ parameters, for varying values of sample size $n$. The conditional distribution is not $\mathrm{SO}(3)$-equivariant in all settings shown.
E.7 Permutation equivariance
@@
In this setting, a convenient maximal invariant to condition on is the order statistic $M(X)=(X_{(1)},\dotsc,X_{(d)})$. For $d\in\{3,4,5\}$ and various values of $s$, the test rejection rates are shown in Figure 14. The results are similar to those of the other synthetic experiments, though we note that the powers of the tests rapidly decrease with increasing dimensions in this particular setting.
Figure 14: Rejection rates for tests for the permutation equivariance experiment for varying values of sample size $n$. The conditional distribution is $\mathbb{S}_{d}$-equivariant only at $s=0$. (Note the different vertical scale for $s=0$ versus $s>0$.)
E.8 Equivariance as joint invariance in the LHC dijet data
@@
Figure 15 shows the distribution of transverse momenta for a sample of $500$ first and second leading constituents from the LHC dijet data. The momenta $X$ of the first leading constituents appear to be distributed on a ring centered at the origin, suggesting that the distribution of $X$ may be marginally invariant with respect to the action of $\mathrm{SO}(2)$. As per Property E7, equivariance could then be tested via a test for joint invariance under the action of the group of paired rotations $\mathbf{G}_{0}=\{(g_{1},g_{2})\in\mathrm{SO}(2)\times\mathrm{SO}(2):g_{1}=g_{2}\}$. We conduct a test for invariance with respect to this subgroup on samples of size $n=25$. Furthermore, as the presence of unexpected symmetry would also be of scientific interest, we also test for invariance under the action of the full $\mathbf{G}_{1}=\mathrm{SO}(2)\times\mathrm{SO}(2)$ group and of $\mathbf{G}_{2}=\mathrm{SO}(4)$. The results based on $B=1000$ randomizations are shown in Table 4. We see that our tests detect $\mathbf{G}_{0}$-invariance and reject $\mathbf{G}_{1}$- and $\mathbf{G}_{2}$-invariance, which is again consistent with the Standard Model.
Test
Group to test BS-FS BS-SK RN-FS RN-SK
$\text{Paired }\mathrm{SO}(2)\text{-rotations}$ $0.044$ $0.052$ $0.050$ $0.052$
$\text{Unpaired }\mathrm{SO}(2)\text{-rotations}$ $0.481$ $0.874$ $0.958$ $0.972$
$\mathrm{SO}(4)$ $0.633$ $0.933$ $0.989$ $0.987$
E.9 Training an invariant MNIST classifier
As an application under a different context, we illustrate how our tests for symmetry of a conditional distribution $P_{Y|X}$ can be adapted as tests for symmetry of a function $f:\mathbf{X}\to\mathbf{Y}$ by taking $P_{Y|X}=\delta_{f(X)}$. This means that our tests could be used as, for example, validation tools for assessing whether a prediction model has learned symmetry after training, requiring only input data and model-predicted output. To demonstrate this utility, consider a setting where an MNIST classifier is trained to be invariant with respect to image rotations via data augmentation (Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022). Unlike methods that encode symmetry as inductive biases, data augmentation may not always succeed in producing a symmetric model.
The MNIST dataset is a collection of $28\times 28$ pixel images of handwritten digits. The training set consists of $20000$ images; the test set consists of $45000$ augmented images, from which we uniformly sample to use in our tests. We train two identical LeNet convolutional neural networks for predicting the digit in the image, where one model is trained on images augmented with rotations and the other without. Before and during training, we pause the training process after every three epochs to test the invariance properties of each model. To generate the data used in each test, a sample of $n=100$ augmented images is sampled from the test set and fed into the model. The data $Y$ are then taken as the digit probabilities predicted by the model, and so each observation is a vector on a simplex of dimension $d=10$ (or $d=9$ in the setting where the “99” images are removed from the data). In our tests, we take the maximal invariant to be the unaugmented image to avoid complications with orienting images.
Figure 16: Test rejection rates for image rotation-invariance of two MNIST models during training, where one is trained with data augmentation and the other without. The top row shows results for when the models are trained on the full MNIST data, and the bottom row shows results for when images with true label 99 are removed from the data.
Here, we used the estimated test rejection rate as a measure of model invariance. It may be tempting to determine whether symmetry has been learned at a particular time point by whether a single test rejects or not, but conducting a suite of tests at multiple points during training would be subject to multiple testing considerations. A potential direction for future work is anytime-valid tests for conditional symmetry, which would allow for the continual evaluation of symmetry-learning based on a single test at each time point during training. Anytime-valid tests for distributional invariance have recently appeared in the literature (Koning, 2023; Ramdas et al., 2023b; Lardy and Pérez-Ortiz, 2025).
@@ E.9 Training an invariant MNIST classifier
[Acknowledgments] The authors would like to thank Arthur Gretton and Antonin Schrab for a helpful conversation about randomization tests and MMD estimators. They are also grateful to the editor, associate editor, and two anonymous referees, whose comments and questions helped improve the paper.